The following population is provided:17, 15, 8, 9, 7, 9, 11, 12, 14, 16, 9, 5, 10, 14, 13, 12, 12, 11, 9, 14, 8, 14, 12. Further, a simple random sample from this population gives the following values: 12, 9, 5, 10, 14, 11. Compute the sampling error for the sample mean.

a. 1.811 b. 1.167 c. 1.181 d. 1.348 e. None of the above


1
Expert's answer
2022-05-10T07:00:30-0400

The sampling error for the sample mean is


"\sigma_{\bar{x}}=\sqrt{\dfrac{\sigma^2}{n}}"

"\mu=\dfrac{1}{23}(17+15+8+9+7+9"

"+11+12+14+16+9+ 5+10"

"+14+13+12+12+11+9+ 14"

"+8+14+12)=\dfrac{261}{23}"

"\sigma^2=\dfrac{\Sigma_i(x_i-\mu)^2}{N}=\dfrac{205.21739130435}{23}"

"\sigma_{\bar{x}}=\sqrt{\dfrac{\sigma^2}{n}}"

"=\sqrt{\dfrac{205.21739130435}{23(6)}}=1.219"

e. None of the above